OpenStudy (oreolover4876):

x^2-12x=4 What number needs to be added to complete the square.

OpenStudy (dangerousjesse):

What was your first impression of this problem? Any guesses?

OpenStudy (oreolover4876):

I honestly have no clue, this is from a section i havent learned yet and my teacher said that a problem like it is going to be on my final

OpenStudy (dangerousjesse):

Okay, I hope you don't mind me taking a minute to explain this haha.

OpenStudy (dangerousjesse):

Complete the square: \[x^{2}-12 x = 4\] Subtract 4 from both sides: \[x^{2}-12 x-4 = 0\] Add 4 to both sides: x^2-12 x = 4 Group terms together on the left hand side, leaving a blank spot to add a constant in later: \[(x^2-12 x+ ) = 4\] Take one half of the coefficient of x and square it, then add it to both sides. Add \[\frac{-12}{2}^{2} = 36\] to both sides: \[(x^{2}-12 x+36) = 4+36\] \[4+36 = 40:\] \[(x^2-12 x+36) = 40\] \[x^2-12 x+36 = (x-6)^2:\]

OpenStudy (oreolover4876):

Thank you sooo much!!

OpenStudy (dangerousjesse):

\[(x-6)^{2}-40=0\]

OpenStudy (dangerousjesse):

No problem :)